OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3)

OpenStudy (shaik0124):

g(3)=3*3+2 =9+2 =11

OpenStudy (anonymous):

ooh okay thank you its becasue i didnt get it thanks:)

OpenStudy (shaik0124):

just u put 3 in x place and solve then u will get answer

OpenStudy (anonymous):

okay do you thinnk you could help me with this -18 = -3 - b please

OpenStudy (shaik0124):

-18+3=-3+3-b -15=-b b=15

OpenStudy (shaik0124):

@sandyhuerta